OpenStudy (thephysicsman):

What is 18 divided by 2 x 3 + 4

OpenStudy (thephysicsman):

When I put it into a calculator it gives me 31 when I do it I get a different answer

OpenStudy (thephysicsman):

I ended up getting 7

OpenStudy (shawn):

screw PEMDAS. Just evaluate from left to right. The answer is what you want it to be

OpenStudy (thephysicsman):

I don't get it isn't that the method you're supposed to use?

OpenStudy (thephysicsman):

Yeah i evaluated it left to right and got 31

OpenStudy (shawn):

Sure. 31 it is.

OpenStudy (thephysicsman):

@mathmale

OpenStudy (anthonyym):

PEMDAS Parentheses Exponents Multiplication Division Addition Subtraction This is in the order of first to last with parentheses first. Since multiplication is before subtraction, you multiply first. So do 2*3 first then add 4 to the product.

geerky42 (geerky42):

"What is *18 divided* by 2 x 3 + 4"

geerky42 (geerky42):

Do multiplication or division first. Order doesn't matter. After that, add 4 to the result.

OpenStudy (mathiscancer):

It's 7

geerky42 (geerky42):

It doesn't matter what order you do in multiplication and division because dividing anything by n is same as multiplying it by 1/n.

OpenStudy (anthonyym):

Is it \[\frac{ 18 }{ 2*3+4 }\] or \[\frac{ 18 }{ 2*3 }+4\]

geerky42 (geerky42):

probably jus18 / 2 * 3 + 4, as in \(\dfrac{18}2\times3+4\)

OpenStudy (mathiscancer):

It's either 7 or 31 depending on how you look at it.

OpenStudy (thephysicsman):

\[18 \div 2*3+4\]

OpenStudy (thephysicsman):

what I did was I multiplied 2*3 first and got 6 then I divided that by 18 and got 3 then added 4 which was 7

OpenStudy (thephysicsman):

@MathIsCancer when I just put the whole thing into the calculator I get 31 but I'm not sure how you would get two different answers

OpenStudy (mathiscancer):

Depends on how you read it

OpenStudy (mathiscancer):

31 makes the most sense, I would go with that.

OpenStudy (thephysicsman):

wait ... I think i see it now.. \[\frac{ 18 }{ 2 }*3+4 = \frac{ 18 }{ 2 }*(\frac{ 3 }{ 1 }) = \frac{ 54 }{ 2 } = 27+4.. = 31..\] thanks

OpenStudy (mathiscancer):

Yup, so 31 is your answer!